f(t) = [5+ (4e2+ + 6te-2)e-24]u(t) A Find F(S) = L[f(t) as a proper rational function. F(s) must be in fully factored form.
Answer #1

as Güven f(t) = (5+ (yezt; stözt) ezt Jult) = [5+4 6+6+ e 4t] uct) =7f(t) = (9+63 4t J ult) Taking the L.T on b.s, we get [{f- +6 etape F(S= } + is this